hartnn (hartnn):

Check my answer! Laplace transform of Integral 0 to t of (1-e^au)/u du i got (1/s) ln [(s-a)/s]

hartnn (hartnn):

\(\Large L[\int \limits_0^t \dfrac{1-e^{au}}{u}du ] = \dfrac{1}{s}\ln \dfrac{s-a}{s}\) is this correct ? can be simplified more?

hartnn (hartnn):

@amistre64 @ganeshie8

OpenStudy (anonymous):

there is this web site mathway it can help u @hartnn

hartnn (hartnn):

thank you :) i know about mathway but i like OpenStudy better :)

hartnn (hartnn):

amistre, if you want, i can upload my work...

OpenStudy (anonymous):

ok sorry ya i like openstudy better i don't blame u LOL

OpenStudy (amistre64):

yeah, uploading your work may help out alot :) \[\Large L[\int \limits_0^t \frac 1u-\frac{e^{au}}{u}du ] \] \[\Large L[\int \limits_0^t \frac {du}u]-L[\int \limits_0^t\frac{e^{au}}{u}du ] \] \[\Large L[ln(t)-ln(0)]-L[e^a\int \limits_0^t\frac{e^{u}}{u}du ] \]

hartnn (hartnn):

i sorta did it some other way...

OpenStudy (amistre64):

some other way is fine ... my way isnt getting me anywhere special lol

OpenStudy (anonymous):

LOL

OpenStudy (amistre64):

you seem to be applying thrms that im just to rusty to verify :/ but the good news is that there are plenty of people smarter than me about :) ... basically anyone with a smartscore above 10 as i see it ;)

OpenStudy (anonymous):

@hartnn, I'm getting the same

hartnn (hartnn):

Thanks :) quick question : did i start the 2nd one correctly ? any better method?

OpenStudy (anonymous):

Just checked your work - I had a slightly different method using a trig sub somewhere along the way, but that's only because I'm not as familiar with the theorems as you are :)

OpenStudy (anonymous):

As for the second one, give me a sec... It doesn't look finished to me.

hartnn (hartnn):

yup, i am solving it right now, as we speak will give you my answer in a minute...

hartnn (hartnn):

my answer for 2nd looks ugly :(

OpenStudy (anonymous):

I'm getting something much messier... \[\begin{align*}\mathcal{L}\left\{t^2\sinh^22t\right\}&=\frac{1}{4}\mathcal{L}\left\{t^2e^{4t}\right\}-\frac{1}{2}\mathcal{L}\left\{t^2\right\}+\frac{1}{4}\mathcal{L}\left\{t^2e^{-4t}\right\}\\ &=\frac{1}{4}\cdot\frac{2!}{(s-4)^3}-\frac{1}{2}\cdot\frac{2!}{s^3}+\frac{1}{4}\cdot\frac{2!}{(s+4)^3}\\ &=\frac{1}{2(s-4)^3}-\frac{1}{s^3}+\frac{1}{2(s+4)^3} \end{align*}\] It's messy if you try to write it as one fraction.

hartnn (hartnn):

oh shoot! i just differentiated once :P

hartnn (hartnn):

yessss now i am getting the same result! :) thank you vey much, i appreciate it :)

OpenStudy (anonymous):

you're welcome!